# Artificial-Intelligence Generated Code Considered Harmful: A Road Map for Secure and High-Quality Code Generation

### v1

@@ 7.3 Code Improvement in LLM
As we discussed in Section 2.2, RL has been employed by LLMs to improve code generation. Due to the implementation of these blackbox components, we would not know for sure if the improvements have been made in the security aspect of code generation. However, our study has shown that LLM-generated code produces a significant amount of security risks. Prompt-based learning [33] is an approach where LLMs are guided to generate desired output with carefully designed prompts without re-training the models. This technique could be applied on code generation as well. For instance, whenever malloc appears in code generation, it should come with an allocation size check before using malloc and a NULL check for the result of malloc. This would be one of the desired outputs to drastically reduce malloc overflow issues. However, based on the results we have seen so far in our study, LLM-generated code still has a long way to go to achieve this level of security.
@@ 8.1 Multiple LLMs
Our study focuses on a single LLM, GPT-4o. We could extend our study by conducting the comparisons using the same human-written code with other LLMs such as Llama3, Claude, etc.; which would reveal differences between different LLMs in terms of code security. However, the focal point of our study is to bring attention to the differences between LLM-generated and human-written code. We believe that GPT-4o, being the state-of-the-art of OpenAI’s LLM, has successfully revealed the security threats brought upon by LLMs-based code generation; the issues we found are critical, yet likely overlooked when programming with the help of LLMs.
@@ 8.2 Coding Tasks
The coding tasks in our experiments are relatively small and independent as opposed to the code in large software projects such as an operating system kernel, a browser etc. Comparing code implemented for a subsystem can be an interesting future work since it requires more attention to interoperability, as these software has a plethora of components within themselves. Another possible future work in terms of coding tasks could be asking LLM to fix known errors or identify possible issues in human-written code.

### v2

@@ 7.3 Code Improvement in LLM
As we discussed in Section 2.2, RL has been employed by LLMs to improve code generation. Due to the implementation of these blackbox components, we would not know for sure if the improvements have been made in the security aspect of code generation. However, our study has shown that LLM-generated code produces a significant amount of security risks. Prompt-based learning [33] is an approach where LLMs are guided to generate desired output with carefully designed prompts without re-training the models. This technique could be applied to code generation as well. For instance, whenever malloc appears in code generation, it should come with an allocation size check before using malloc and a NULL check for the result of malloc. This would be one of the desired outputs to drastically reduce malloc overflow issues. However, based on the results we have seen so far in our study, LLM-generated code still has a long way to go to achieve this level of security.
@@ 8.1 Multiple LLMs
Our study focuses on a single LLM, GPT-4o. We could extend our study by conducting the comparisons using the same human-written code with other LLMs such as Llama3, Claude, etc.; which would reveal differences between different LLMs in terms of code security. However, the focal point of our study is to bring attention to the differences between LLM-generated and human-written code. We believe that the state-of-the-art LLM at the time of writing, GPT-4o, has successfully revealed the security threats brought upon by LLMs-based code generation; the issues we found are critical, yet likely overlooked when programming with the help of LLMs.
@@ 8.2 Coding Tasks
The coding tasks in our experiments are relatively small and independent as opposed to the code in large software projects such as an operating system kernel, a browser, etc. Comparing code implemented for a subsystem can be an interesting future work since it requires more attention to interoperability, as these software has a plethora of components within themselves. Another possible future work in terms of coding tasks could be asking LLM to fix known errors or identify possible issues in human-written code.